Lucifer February Com Night 5 February 10-11 2009

Observers: N.Ageorges, W.Seifert, V.Knierem (LUCIFER), DThompson (@LBT), JHill (Tucson)
Telescope Operator: W.Wack
SW Support: TSargent
Telescope Support: DOfficer, JMorris

Summary:

Open at sunset with fairly thick clouds but it does not look like it will rain.

22:19 MST Tout=-3.9 Dewpt=-13.1 RH=45.8% and rising Wind=10-15m/s

We had an instance of the GCS-AGW communication problem earlier. TSargent noted the time it happened. It looked like the auto-restart was trying, but repeatedly failing. We restarted GCSL but the problem persisted. AGW was stopped, restarted, and the stages re-homed.

Stuck mask on LUCIFER (sometime before 23:00). One of the long-slit masks got stuck in the focal plane. 04:30 fixed, mask back in storage unit.

23:02 Closing, heavy clouds and some structure on the radar and satellite loops.

04:45 Tout=-3.5C, RH=46.3%, Dewpt=-13.4 wind now gusting to 23m/s.

GCSGUI Suggestion from John: When GCS is paused (shown on the IIF GUI) and then a large offset comes through that would move the guide probe out of the patrol field, GCS shows an ERROR and the background is red. GCS should know that it was paused and thus should show this as a warning.

-- DavidThompson - 12 Feb 2009
Topic revision: r2 - 12 Feb 2009, DavidThompson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback